Unit 10 - Financial Statement Analysis

Following is a comparative balance sheet for Summer Days Corporation:

                                                                        

                                                Summer Days Corporation

                                                Comparative Balance Sheet

                                               December 31, 2019 and 2018

                                                                                                2019                      2018

Current assets:

Cash                                                                               $ 140,000               $ 90,000

Short-term investments                                                  90,000                  80,000

Accounts receivable, net                                              350,000                220,000

Inventory                                                                         500,000                430,000

Prepaid expenses                                                             30,000                  30,000

Total current assets                                                  $1,110,000             $ 850,000

Property, plant, and equipment, net                         750,000                500,000

Other assets                                                                     280,000                300,000

Total assets                                                                 $2,140,000           $1,650,000

Current liabilities:

Short-term notes payable                                           $650,000              $670,000

Accounts payable                                                          200,000                225,000

Total current liabilities                                               $850,000              $895,000

Non-current debt                                                           380,000                220,000

Total liabilities                                                           $1,230,000           $1,115,000

Shareholders' equity:

Common shares                                                           $500,000              $500,000

Retained earnings                                                          410,000                  35,000

Total shareholders' equity                                        $ 910,000              $535,000

Total liabilities and shareholders' equity            $2,140,000           $1,650,000

Notes: Net sales (all on credit) and cost of goods sold for the year ended December 31, 2019, were $2,000,000 and $1,200,000, respectively. The number of common shares outstanding has been 50,000 since the company began operations.

Required:

Calculate the following ratios for the year ended December 31, 2019:

            a. current ratio

            b. acid-test ratio

            c. inventory turnover

            d. accounts receivable turnover

            e. debt ratio

            f. book value per common share of stock
Answer #1
a.
Current ratio = Total current assets / Total current liabilities = 1110000 / 850000 = 1.31
b.
Acid-test ratio = ( Total current assets - Inventory - Prepaid expenses ) / Total current liabilities = ( 1110000 - 500000 - 30000 ) / 850000 = 0.68
c.
Average inventory = (beginning inventory + Ending inventory) / 2 = (430000 + 500000) / 2 = 465000
Inventory turnover = Cost of goods sold / Average inventory = 1200000 / 465000 = 2.58
d.
Average accounts receivable = (beginning accounts receivable + Ending accounts receivable ) / 2 = ( 220000 + 350000 ) / 2 = 285000
Accounts receivable turnover = Net Credit Sales / Average accounts receivable = 2000000 / 285000 = 7.02
e.
Debt ratio = Total liabilities / Total assets = 1230000 / 2140000 = 57.48%
f.
Book value per common share of stock = Total stockholders equity / Shares of common stock outstanding = 910000 / 50000 = 18.20
Note : All the answers are rounded off to 2 decimal places
If you have any doubt then please ask